# Multiview learning with twin parametric margin SVM

### v1

@@ 4.5 Effect of hyperparameters $C_{1}$ and $\sigma$
In this subsection, we conduct an in-depth investigation to understand the influence of the hyperparameters $C_{1}$ and $\sigma$ on the comprehensive predictive capacity of the proposed MvTPMSVM model. Figure 7 depicts the results of sensitivity analysis conducted on both KEEL and UCI datasets. The accuracy is assessed by altering the parameters $C_{1}$ and $\sigma$. It is observable that with the escalation of $C_{1}$ and $\sigma$ values, there is a corresponding increase in accuracy. However, once a specific threshold is exceeded, the accuracy stabilizes, indicating that further increments in $C_{1}$ and $\sigma$ exhibit a diminishing effect on testing accuracy. Consequently, meticulous selection of hyperparameters for the proposed MvTPMSVM model is imperative to attain optimal generalization performance.
@@ 5 Conclusions and Future Work
In this paper, we proposed a novel multiview twin parametric margin support vector machine (MvTPMSVM) to address the challenges faced by MvTSVM. The proposed MvTPMSVM model alleviates the influence of heteroscedastic noise present in the data, simultaneously reducing the computational costs typically associated with SVM-2K, MvTSVM, and their variants. Furthermore, the proposed MvTPMSVM model directly applies the kernel trick for nonlinear cases, enabling it to solve the exact formulation. In order to showcase the effectiveness and efficiency of the proposed MvTPMSVM model, we performed an extensive series of experiments and subjected them to thorough statistical analyses. We conducted experiments on $55$ UCI and KEEL datasets, and these experimental results underwent a comprehensive statistical assessment employing a ranking scheme, Friedman test, Nemenyi post-hoc test, and win-tie-loss sign test. Our findings from the experiments, combined with the statistical analyses, indicate that the proposed MvTPMSVM model outperforms baseline models in terms of generalization performance. Furthermore, the proposed model underwent testing on both synthetic and $45$ real-world AwA datasets. The outcomes demonstrated notable enhancements in accuracy and generalization performance by employing multiple views in the analysis. The proposed MvTPMSVM model exhibits superior performance compared to certain early and late fusion models, as well as selected state-of-the-art multiview models. Our research has notable limitations, such as focusing on binary classification and utilizing only two views. In the future, we aim to expand this framework to encompass multiple classifications and incorporate more than two views for a more comprehensive analysis. While our proposed models have shown exceptional performance in binary classification problems, their evaluation in multiview (incorporating more than two views) multiclass problems has not been conducted. A critical avenue for future research would involve adapting these models to be suitable for multiview multiclass problems. Also, our model designed to handle balanced datasets, has the potential to enhance its robustness by integrating techniques tailored to address class imbalance learning. This refers to scenarios where the distribution of classes in the training data is highly skewed, with one or more classes significantly outnumbering others. By incorporating methods specifically designed to tackle class imbalance, the model becomes more adept at effectively handling datasets with disparate class distributions. The source code of the proposed MvTPMSVM model is available at https://github.com/mtanveer1/MvTPMSVM.

### v2

@@ 4.5 Effect of hyperparameters $C_{1}$ and $\sigma$
In this subsection, we conduct an in-depth investigation to understand the influence of the hyperparameters $C_{1}$ and $\sigma$ on the comprehensive predictive capacity of the proposed MvTPMSVM model. Figure 7 depicts the results of sensitivity analysis conducted on both UCI and KEEL datasets. The ACC is assessed by altering the parameters $C_{1}$ and $\sigma$. It is observable that with the escalation of $C_{1}$ and $\sigma$ values, there is a corresponding increase in ACC. However, once a specific threshold is exceeded, the ACC stabilizes, indicating that further increments in $C_{1}$ and $\sigma$ exhibit a diminishing effect on testing ACC. Consequently, meticulous selection of hyperparameters for the proposed MvTPMSVM model is imperative to attain optimal generalization performance.
@@ 5 Conclusions and Future Work
In this paper, we proposed a novel multiview twin parametric margin support vector machine (MvTPMSVM) to address the challenges faced by MvTSVM. The proposed MvTPMSVM model alleviates the influence of heteroscedastic noise present in the data, simultaneously reducing the computational costs typically associated with SVM-2K, MvTSVM, and their variants. Furthermore, the proposed MvTPMSVM model directly applies the kernel trick for nonlinear cases, enabling it to solve the exact formulation. In order to showcase the effectiveness and efficiency of the proposed MvTPMSVM model, we performed an extensive series of experiments and subjected them to thorough statistical analyses. We conducted experiments on $55$ UCI and KEEL datasets, and these experimental results underwent a comprehensive statistical assessment employing a ranking scheme, Friedman test, Nemenyi post-hoc test, and win-tie-loss sign test. Our findings from the experiments, combined with the statistical analyses, indicate that the proposed MvTPMSVM model outperforms baseline models in terms of generalization performance. Furthermore, the proposed model underwent testing on both synthetic and $45$ real-world AwA datasets. The outcomes demonstrated notable enhancements in ACC and generalization performance by employing multiple views in the analysis. The proposed MvTPMSVM model exhibits superior performance compared to certain early and late fusion models, as well as selected state-of-the-art multiview models. Our proposed model has shown exceptional performance in binary classification problems, their evaluation in multiview (incorporating more than two views) multiclass problems has not been conducted. A critical avenue for future research would involve adapting the proposed model to be suitable for multiview multiclass problems. Also, our model designed to handle balanced datasets, has the potential to enhance its robustness by integrating techniques tailored to address class imbalance learning. This refers to scenarios where the distribution of classes in the training data is highly skewed, with one or more classes significantly outnumbering others. By incorporating methods specifically designed to tackle class imbalance, the model becomes more adept at effectively handling datasets with disparate class distributions. The source code of the proposed MvTPMSVM model is available at https://github.com/mtanveer1/MvTPMSVM.